OpenStudy (anonymous):

two planes leave the airport at the same time , one flying at 300km/h and the other at 420 km.h. The angle between their flight paths is 90 degrees. after 3 hours how far apart are they?

OpenStudy (anonymous):

\[\sqrt{(3*300)^2+(3*420)^2}=180 \sqrt{74}=1548.42 \text{km} \]